You are here: Foswiki>FLAO Web>LbtLab>LabLog20100220 (21 Feb 2010, EnricoPinna)Edit Attach

Alfio, Carmelo, Simone, Enrico, Mauro, Paolo, Luca

We found that alignment tests have been perfomred at a wrong value of modulation amplitude of the TT mirror.
Up to now we always tested the system using 8V @100Hz while in Arcetri the usual modulation for the alignment check was 6V@100Hz.

TT MALFUNCTION (more illumination of two pupils opposite to the center)
Checking the images of the TTcalibration we found that at 8V the TT mirrror can no longer produce a circular pattern in the focal plane and the circle is elongated. This effect can explain the malfunction detected yasterday. We reproduced the same malfunction rotating the AGW of 90deg and 180deg usign 8V@100Hz, while disappear at 6V100Hz.

PUPIL SQUARE SIDE
Moreover at 8V@200Hzthe side of the square geometry of the pupils is 36.2 instead the 36.0 required and we measure now using the correct modulation value of 6V@100Hz.

LC focusing setup
A reference stop has been placed on the mount of the LC. 2 shim of 90um each has been placed in order to be ready to tranlate LC of 180um.

CCD47 Cleaning
CCD47 has been unmounted and its CCD window has been cleaned

-- EnricoPinna - 20 Feb 2010
Topic revision: r3 - 21 Feb 2010, EnricoPinna
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback